OpenStudy (anonymous):

In 1992, the life expectancy of males in a certain country was 67.4 years. in 1998, it was 69.7 years. Let E represent the life expectancy in year t and let t represent the numbr of years since 1992. Q; The linear function that fits the data is E(t)=?t+? (round to the nearest tenth) Q; Use the function to predict the life expectancy pf males in 2004 E(12)=?

OpenStudy (anonymous):

69.7 - 67.4 = 2.3 slope = 2.3/6 = 0.3833 function is E = 0.3833t + 67.4 in 2004 t = 12 E = 0.38333*12 + 67.4 = 72

OpenStudy (anonymous):

is that rounded to the nearest tenth?

OpenStudy (anonymous):

yea - it came out to 71.99

OpenStudy (anonymous):

AWESOME! thanks a lot!

OpenStudy (anonymous):

no probs

OpenStudy (anonymous):

plz check good answer to close this question

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

ty